The area of a rectangle is 24,396 square centimeters. If the width is 38 centimeters what is the length? How do you know?
Answer by Alan3354(69443)

The area of a rectangle is 24,396 square centimeters. If the width is 38 centimeters what is the length?
Area = L*W
L = Area/W = 24396/38
L = 642 cm
--------------
How do you know?
My dog told me.